Four unequal resistors connected in series have same current but different voltages. O True False
Answer #1

Yes it is true.

When resistors are connected in series the current through the circuit is same in all resistors.

Current (I) = Voltage /resultant resistant

Resultant resistantance(r) =r1+r2+r3+r4

I =V/r

Voltage in each resistor is

V1=i×r1

V2=i×r2

V3=i×r3

V4=i×r4